# Policy Space Response Oracles: A Survey

### v1

@@ 8 Applications
Game-theoretic analysis in PSRO relies on game models, which abstract away the underlying game structures. This abstraction enables PSRO to solve a variety of games or address issues that can be formulated as a game, and yields numerous applications of PSRO in disparate domains. Specifically, PSRO has been applied to specialized games, including security games Wang et al. (2019); Wright et al. (2019); Fang (2019); Tong et al. (2020); Xu et al. (2021); Cui and Yang (2023), bargaining games Li et al. (2023b); Wang and Wellman (2024), Colonel Blotto games An and Zhou (2023), Pursuit-Evasion games Li et al. (2023a), auctions Li and Wellman (2021), and mechanism design Zhang et al. (2023).
@@
Besides the real-world applications, PSRO (including DO) has also been applied for designing novel algorithms in domains that can be modeled as a game. For example, the PSRO framework has facilitated developments in robust policy discovery within robust reinforcement learning domains Liang et al. (2023), value alignment in large language models Ma et al. (2023), public health services Killian et al. (2023), and the discovery of information in images Giboulot et al. (2023).
In the field of computer vision, several contributions towards data augmentation for object detection Behpour et al. (2019a), active learning Behpour et al. (2019b), semi-supervised multi-label classification Behpour (2018), and video tracking Fathony et al. (2018) underscore the PSRO framework’s impact on enhancing computer vision and structured prediction methodologies. Furthermore, PSRO-type methods have been adapted to enhance the training of Generative Adversarial Networks (GAN) Oliehoek et al. (2019); Aung et al. (2022), with notable implementations such as DO-GAN, which facilitates GAN training in high-dimensional spaces through strategy pruning and continual learning mechanisms.
@@ 9 Open Research Questions
In the standard PSRO framework, an empirical game is represented in normal form. As the number of players increases, the normal-form representation expands exponentially, leading to a significant increase in the cost of evaluating the empirical game. While this problem can be mitigated in certain special game types such as symmetric games, the practicality of applying PSRO becomes impractical when dealing with a very large number of players. A potential remedy involves employing game model learning Sokota et al. (2019), essentially a utility function regression. The acquired utility function aims to extrapolate across the (empirical) strategy space, eliminating the need to evaluate each individual profile.
Another possible approach is to use polymatrix game representations. A polymatrix game is a many-player game where each player can be thought of as a node in an interaction graph and on every edge in this graph there is a bimatrix game that the node plays with its neighbours, gaining as utility the sum of payoffs across these bimatrix games. Polymatrix games have been used within equilibrium computation methods for many-player normal-form games, due to their linear formulation. In the work of both Govindan and Wilson Govindan and Wilson (2004) and Gemp at al. Gemp et al. (2022) a polymatrix approximation of the input game is maintained and updated (implicitly in the latter case). There appears to be potential to combine or extract elements from these methods to build polymatrix-based empirical game models for PSRO-type methods.
@@
Both best-response computation and empirical game simulation can benefit from parallelization. Firstly, best-response computations among players are parallelizable as they are independent. Secondly, the evaluation of strategy profiles in the empirical game can be parallelized. Additionally, since the evaluation of each strategy profile involves running simulations multiple times and averaging the results, these simulation runs can also be parallelized. Consequently, the PSRO framework lends itself well to high parallelization. Despite this potential, there is currently no PSRO implementation that fully harnesses these three aspects.
@@
Other game-theoretic RL approaches, such as those based on counterfactual regret minimization (CFR) Brown et al. (2019); McAleer et al. (2023a) or policy gradients Fu et al. (2022); Perolat et al. (2022), could be potentially combined with PSRO. In contrast to PSRO, these approaches solve the entire game upfront, and as a result may be slower to converge on certain games and faster on others. CFR-based and policy-gradient-based methods might perform better in games where mixing carefully at many decision nodes is crucial, because PSRO usually mixes only at the root of the game and as a result might require many policies to effectively mix at information sets McAleer et al. (2021). Notably, XDO does mix at every decision node, and as a result is a great candidate for hybrid approaches with CFR-based and policy-gradient-based techniques. In particular, it would be useful to automatically decide on which decision nodes to mix. Since some games have sections that do not require mixing such as complex control tasks, those would be more efficient to use the deep RL policy from PSRO or XDO, while nodes that require careful mixing should defer to a CFR-based or policy-gradient-based technique.
@@
Hyperparameters exist universally in PSRO variants, including the lower bound of the probability of playing a strategy in PRD Lanctot et al. (2017), the regret threshold in RRD Wang and Wellman (2023b), the probability of employing the uniform MSS Wang et al. (2019), and the weights for diversity regularizers Perez-Nieves et al. (2021). These hyperparameters often require deliberate tuning to fit a specific game by running the same experiments multiple times for different hyperparameter choices, which can be costly. Although some of the hyperparameters can be set heuristically (e.g., by employing a simple annealing scheme), the question of how to tune the hyperparameters automatically and efficiently remains open. Specifically, can we design an automatic adjustment scheme for these hyperparameters such that PSRO can be directly applied to any game without manual game-specific fine turning?
@@
Acknowledgement
We thank Michael P. Wellman for his invaluable advice on improving the accuracy and readability of the survey.

### v2

@@ 8 Applications
Game-theoretic analysis in PSRO relies on restricted games, which abstract away the underlying game structures. This abstraction enables PSRO to solve a variety of games or address issues that can be formulated as a game, and yields numerous applications of PSRO in disparate domains. Specifically, PSRO has been applied to specialized games, including security games Wang et al. (2019); Wright et al. (2019); Fang (2019); Tong et al. (2020); Xu et al. (2021); Cui and Yang (2023), bargaining games Li et al. (2023c); Wang and Wellman (2024), Colonel Blotto games An and Zhou (2023), Google Research Football environments Liu et al. (2021); Song et al. (2024), chess Zahavy et al. (2023); Li et al. (2023b), Pursuit-Evasion games Li et al. (2023a, 2024a), auctions Li and Wellman (2021), and mechanism design for sequential auctions Zhang et al. (2023), which are among a class of the hardest extensive-form games to solve.
@@
Besides these real-world applications, PSRO (including DO) has also been applied for designing novel algorithms in domains that can be modeled as a game. For example, the PSRO framework has facilitated developments in RL for robust policy discovery Liang et al. (2023) and policy generalization Yang et al. (2022), multiagent RL evaluation Li and Wellman (2024), value alignment in large language models Ma et al. (2023), public health services Killian et al. (2023), combinatorial optimization Wang et al. (2024), and the discovery of information in images Giboulot et al. (2023).
PSRO has also had an impact on enhancing computer vision and structured prediction methodologies, for example, via data augmentation for object detection Behpour et al. (2019a), active learning Behpour et al. (2019b), semi-supervised multi-label classification Behpour (2018), and video tracking Fathony et al. (2018). Furthermore, PSRO-type methods have been adapted to enhance the training of Generative Adversarial Networks (GAN) Oliehoek et al. (2019); Aung et al. (2022).
9 PSRO Implementations
Two software libraries that include PSRO implementations are OpenSpiel Lanctot et al. (2019) and MALib Zhou et al. (2023). Both serve as comprehensive toolsets including various games and algorithms for exploring general reinforcement learning and search or planning in games.
@@ 10 Open Research Questions
In the standard PSRO framework, an empirical game is represented in normal form. As the number of players increases, the normal-form representation expands exponentially, leading to a significant increase in the cost of evaluating the empirical game. While this problem can be mitigated in certain special game types such as symmetric games, the practicality of applying PSRO diminishes when dealing with a very large number of players. A potential remedy involves employing game model learning Sokota et al. (2019), essentially a utility function regression. The acquired utility function aims to extrapolate across the (empirical) strategy space, eliminating the need to evaluate each individual profile.
Another possible approach is to use polymatrix game representations. A polymatrix game is a many-player game where each player can be represented as a node in an interaction graph. On every edge in this graph, there is a bimatrix game that the node plays with its neighbors, with the utility for each player being the sum of payoffs from these bimatrix games. Polymatrix games have been used within equilibrium computation methods for many-player normal-form games, due to their linear formulation. In the work of both Govindan and Wilson Govindan and Wilson (2004) and Gemp at al. Gemp et al. (2022) a polymatrix approximation of the input game is maintained and updated (implicitly in the latter case). There appears to be potential to combine or extract elements from these methods to build polymatrix-based empirical game models for PSRO-type methods.
@@
Both best-response computation and empirical game simulation can benefit from parallelization. Firstly, best-response computations among players are parallelizable as they are independent. Secondly, the evaluation of strategy profiles in the empirical game can be parallelized. Additionally, since the evaluation of each strategy profile involves running simulations multiple times and averaging the results, these simulation runs can also be parallelized. Consequently, the PSRO framework is highly suitable for parallelization. Despite this potential, there is currently no PSRO implementation that fully harnesses these three aspects.
@@
CFR-based and policy-gradient-based methods might perform better in games where mixing carefully at many decision nodes is crucial, providing more granular mixing at different game scales (e.g., information sets that define subgames). In contrast, PSRO usually mixes only at the root of the game, limited to distributions over strategies in the restricted game, and might require many policies to effectively mix at information sets McAleer et al. (2021). Therefore, for game sections that do not require mixing such as complex control tasks, it would be more efficient to use the deep RL policies from PSRO or XDO, while decision nodes that require careful mixing should defer to a CFR-based or policy-gradient-based technique. Overall, a hierarchical PSRO structure that controls the granularity of the mixing might be needed.
@@
Hyperparameters exist universally in PSRO variants, including the lower bound of the probability of playing a strategy in Projected Replicator Dynamics Lanctot et al. (2017), the regret threshold in RRD Wang and Wellman (2023b), the probability of employing the uniform MSS Wang et al. (2019), and the weights for diversity regularizers Perez-Nieves et al. (2021). These hyperparameters often require deliberate tuning to fit a specific game by running the same experiments multiple times for different hyperparameter choices, which can be costly. Although some of the hyperparameters can be set heuristically (e.g., by employing a simple annealing scheme), the question of how to tune the hyperparameters automatically and efficiently remains open. Specifically, can we design an automatic adjustment scheme for these hyperparameters such that PSRO can be directly applied to any game without manual game-specific fine turning? As a recent attempt, Li et al. Li et al. (2024b) proposed to learn a policy for hyperparameter optimization that can automatically select the optimal hyperparameter values in PSRO during game solving. Moreover, they highlighted that a well-trained policy is generalizable to different games, thus reducing the cost of hyperparameter tuning when applying PSRO to various games.
@@
Acknowledgements
We thank Marc Lanctot, Karl Tuyls, Michael P. Wellman, Yaodong Yang, and their students, for their invaluable advice on improving the survey.